OpenStudy (electrochika):

dy/dx of y= sin^2(3x-2)

jimthompson5910 (jim_thompson5910):

\[\Large y = \sin^2(3x-2)\] \[\Large \frac{dy}{dx} = 2\sin(3x-2)\cdot\frac{d}{dx}(\sin(3x-2))\cdot\frac{d}{dx}(3x-2)\] \[\Large \frac{dy}{dx} = 2\sin(3x-2)\cdot\cos(3x-2)\cdot(3)\] \[\Large \frac{dy}{dx} = 6\sin(3x-2)\cos(3x-2)\]

OpenStudy (electrochika):

Thank you!!

jimthompson5910 (jim_thompson5910):

yw